How many grams of ammonia will be produced if there are 500.0 moles of nitrogen gas?

1
Expert's answer
2021-11-12T13:01:01-0500

N2(g) + 3H2-> 2NH3(g) This is the balanced equation

Molar mass of N2= 17.031

500/28.03= 17.81

17.81/2=8.905

Molar mass of NH3= 17.031

= 8.905×17.031= 151.66g